solve or simplify this sqrt (x-11)^2+(y-11)^2-sqrt(x+11)^2+(y+11)^2=22

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question

(x-11)^2 + (y-11)^2 - \sqrt{(x+11)^2 + (y+11)^2} = 22
x^2 + y^2 - 22x - 22y + 242 - \sqrt{x^2+y^2+22x+22y+242} = 22
x^2 + y^2 - 22x - 22y - \sqrt{x^2+y^2+22x+22y+242} = -220
x^2 - \sqrt{x^2+y^2+22x+22y+242} - 22x + y^2 - 22y + 242 = 22